# DiagECG: An LLM-Driven Framework for Diagnostic Reasoning via Discretized ECG Tokenization

### v1

@@ Abstract
Electrocardiography (ECG) plays a central role in cardiovascular diagnostics, yet existing automated approaches often struggle to generalize across clinical tasks and offer limited support for open-ended reasoning. We present DiagECG, a novel framework that integrates time-series (TS) and language modeling by enabling large language models (LLMs) to process 12-lead ECG signals for clinical text generation tasks. Our approach discretizes continuous ECG embeddings into symbolic tokens using a lead-independent encoder and quantization module. These symbolic representations are then used to extend the LLM’s vocabulary with ECG-specific tokens, allowing the model to handle both ECG and natural language inputs in a unified manner. To bridge the modality gap, we pretrain the model on an autoregressive ECG forecasting task, enabling the LLM to model temporal dynamics using its native language modeling capabilities. Finally, we perform instruction tuning on both ECG question answering and diagnostic report generation. Without modifying the core model, DiagECG achieves strong performance across tasks while maintaining generalization to out-of-distribution settings. Extensive experiments demonstrate the effectiveness of each component and highlight the potential of integrating symbolic ECG representations into LLMs for medical reasoning.
@@ Introduction
Figure 1: UMAP visualizations of ECG and text embedding distributions across different models. (a) TEST exhibits clear modality separation. (b) TIMELLM produces unstructured and scattered embeddings. (c) CLIP shows partially aligned clusters, but still fragmented modality boundaries. (d) DiagECG achieves compact, overlapping embeddings without contrastive pairing, indicating effective integration of ECG and text in a shared semantic space.
@@
Cardiovascular diseases (CVDs) are the leading cause of mortality and morbidity worldwide (Di Cesare et al. 2023), and electrocardiography (ECG) remains a fundamental tool for early diagnosis and monitoring. With the increasing volume and complexity of ECG data in clinical settings, computer-aided interpretation has become an essential supplement to clinical expertise. Recent advances in deep learning have significantly improved performance in ECG classification tasks (Ge et al. 2024; Guhdar, Mohammed, and Mstafa 2025; Na et al. 2024a; Nejedly et al. 2021; Oh et al. 2022), surpassing traditional signal-processing-based pipelines (Chen et al. 2022a). However, most existing models are tailored to closed-set classification and rely heavily on supervised training. When confronted with new diagnostic categories or unseen clinical tasks, they require extensive retraining with labeled data, limiting their scalability and adaptability. Moreover, these models are typically designed for narrow classification objectives and cannot support more flexible reasoning tasks such as open-ended question answering or narrative report generation.
@@
In parallel, large language models (LLMs) have demonstrated impressive generalization in vision-language domains (Alayrac et al. 2022; Huang et al. 2023), prompting efforts to extend such capabilities to physiological TS data like electrocardiograms (ECGs). However, this extension poses unique challenges: unlike symbolic and semantically structured text, ECG signals are continuous, noisy, and lack explicit semantic boundaries. These fundamental differences complicate the integration of ECG and text into a unified representation space. Recent studies have explored multimodal LLMs for clinical applications, including ECG-based question answering (Liu et al. 2024a; Zhao et al. 2024) and report generation (Wan et al. 2024). Most adopt a two-stage paradigm where an ECG encoder produces continuous representations that are subsequently consumed by a language model. However, this design suffers from two major limitations. First, the modality gap between continuous ECG embeddings and discrete textual semantics remains underexplored. Prior alignment strategies in time series–language modeling (Jin et al. 2024; Sun et al. 2024) or contrastive learning in vision–language settings (Radford et al. 2021) often fail to achieve consistent cross-modal alignment. These limitations, when directly applied to ECG, can lead to unstable integration, as illustrated in Figure 1. Second, ECG signals exhibit high redundancy due to waveform repetition and stable low-variance segments. Directly feeding such continuous features into LLMs can lead to overfitting and reduced generalization to rare or subtle signal variations. These challenges highlight the need for a more structured and symbolic interface between ECG signals and language models.
@@
To address the challenges of bridging physiological signals and language models, we introduce DiagECG, a unified framework that enables LLMs to reason over ECG signals in an open-ended, instruction-driven manner. Instead of relying on paired ECG–text supervision or complex cross-modal alignment losses, DiagECG represents ECGs as discrete symbolic tokens, naturally compatible with the LLM’s vocabulary.
@@
Our method is built on three core contributions: First, we design a lead-wise encoder that processes each ECG lead independently using an identical encoder built from temporal inception enhancement blocks. This architecture captures waveform-specific temporal patterns while avoiding inter-lead interference, resulting in precise and interpretable representations. Second, we introduce a discretization-based tokenizer that converts continuous encoder outputs into symbolic tokens via a fixed-scale quantizer (FSQ) (Mentzer et al. 2023). These tokens are optimized through autoregressive pretraining and embedded in a shared space with language, enabling seamless integration with LLMs without contrastive learning or paired ECG-text data. Third, we apply lightweight instruction tuning for ECG report generation and question answering. By updating only low-rank modules and prompting with structured clinical inputs, our method achieves strong zero-shot generalization across tasks and datasets with minimal supervision.
@@
DiagECG is evaluated on two ECG understanding benchmarks: question answering and diagnostic report generation. It achieves state-of-the-art performance across multiple datasets and exhibits strong zero-shot generalization. Further analysis confirms the contribution of each design component, and visualizations illustrate interpretable attention over clinically meaningful waveform regions.
@@
Recent work has begun to explore the integration of large language models (LLMs) into ECG interpretation tasks, particularly for open-ended question answering and diagnostic report generation. ECG-ReGen (Tang et al. 2025) adopts a retrieval-augmented framework that identifies similar historical cases to assist report generation in a zero-shot setting. PULSE (Liu et al. 2024a) constructs an instruction-tuning dataset based on ECG images, enabling the fine-tuning of image-text multimodal models. MEIT (Wan et al. 2024) introduces TS embeddings directly into the LLM’s attention layers to bridge modality differences, while (Tang et al. 2024) treats ECG features as soft prompts and applies meta-learning for rapid adaptation to new tasks. Despite these advances, most existing methods depend on paired ECG–text data to supervise modality alignment, which requires costly annotation and constrains applicability. In contrast, our approach aligns ECG and language representations through token-level pretraining, avoiding the need for paired supervision while maintaining flexibility across tasks.
@@
Large language models have been increasingly applied to TS tasks, typically via two main directions. One leverages LLMs to extract textual features to guide TS modeling (Wang et al. 2024a; Liu et al. 2025b, a). The other treats LLMs as core TS processors, mainly via two paradigms. The TS-as-text approach (Wang et al. 2025; Ansari et al. 2024; Wang et al. 2024b; Gruver et al. 2023) tokenizes numerical sequences into text for direct input into LLMs, but suffers from length and attention complexity limitations. The TS-embedding-based approach instead encodes signals into continuous embeddings and feeds them into frozen LLMs, often using prompt tuning or normalization (Zhou et al. 2023; Liu et al. 2024c). Some recent work aligns TS embeddings with language semantics, such as S2IP (Pan et al. 2024), TEST (Sun et al. 2024), and TIME-LLM (Jin et al. 2024), through contrastive or reprogramming strategies. Others like VITRO (Bellos, Nguyen, and Corso 2025) and UniTime (Liu et al. 2024b) propose task-specific vocabularies or reconstruction pretraining. However, most focus on forecasting or classification, with limited support for open-ended language tasks grounded in physiological signals. In contrast, our method discretizes ECG features into symbolic tokens, enabling autoregressive pretraining and unified ECG-text processing for clinical QA and report generation.
@@
Figure 2: Overview of the DiagECG framework. The model consists of three stages: (1) ECG Tokenizer, where 12-lead ECG signals are processed by a lead-wise encoder followed by a FSQ module to produce discrete ECG tokens; (2) ECG Token Pretraining, where a language model is trained on the generated tokens with full fine-tuning to capture temporal and morphological patterns; and (3) QA/Report Tuning, where the pretrained model is adapted to downstream tasks using lightweight LoRA tuning. Prompts include structured tabular features and textual instructions to guide generation or question answering.
Given a collection of 12-lead ECG signals $X=\{x_{i}\}_{i=1}^{n}$, where each $x_{i}\in\mathbb{R}^{L\times C}$ consists of $L=5000$ time steps and $C=12$ leads, we consider two tasks: ECG question answering (ECG-QA) and diagnostic report generation (ECG-Report).
@@
In this task, the model generates a diagnostic report $r_{i}\in T$ for each ECG signal $x_{i}$, where $T$ denotes the space of medical text. The goal is to learn a mapping $f_{\mathrm{report}}:x_{i}\mapsto\hat{r}_{i}$ that produces clinically coherent and diagnostically relevant reports.
@@ Methods
We propose DiagECG, a unified framework that enables large language models (LLMs) to perform open-ended clinical reasoning over ECG signals through symbolic representation and lightweight adaptation.
First, we design a lead-wise encoder that captures fine-grained temporal features independently for each ECG lead, enhancing representation fidelity without cross-lead interference. Second, we introduce a discretization mechanism that maps continuous features to symbolic tokens using an FSQ, forming an ECG-specific vocabulary compatible with LLMs. These tokens are pretrained via autoregressive prediction, naturally aligning time series and text without paired supervision. Finally, DiagECG is instruction-tuned on clinical tasks such as ECG report generation and question answering, using only LoRA modules for efficient adaptation. An overview of the full framework is shown in Figure 2.
@@
Inception Block. The Inception Block applies multiple 1D convolutions with kernel sizes $\{1,3,5,7\}$ in parallel to extract multi-scale temporal features. The outputs are summed and passed through GroupNorm and a LeakyReLU activation, followed by a residual connection to preserve information flow:
@@
Temporal Inception Enhancement (TIE) Block. To further refine temporal dependencies, we introduce the TIE Block, which stacks multiple residual convolution layers with downsampling. Each block expands the receptive field while maintaining efficient parameter usage. These blocks are applied sequentially, enabling the encoder to model long-range signal dependencies and compress the temporal resolution:
$\mathbf{h}^{(i+1)}=\mathrm{TIEBlock}(\mathbf{h}^{(i)}),\quad i=1,\dots,N.$
This hierarchical structure allows the encoder to balance fine-grained waveform preservation with contextual abstraction.
@@
To obtain symbolic representations, each latent vector at timestep $t$ $\mathbf{z}_{t}\in\mathbb{R}^{D^{\prime}}$ is projected into a lower-dimensional space $\mathbb{R}^{D}$ using a linear layer with parameters $\mathbf{W}\in\mathbb{R}^{D\times D^{\prime}}$ and $\mathbf{b}\in\mathbb{R}^{D}$:
@@
where $\sigma(\cdot)$ is the sigmoid function. The quantized code is obtained by uniformly quantizing each projected dimension into $L$ discrete levels:
$\mathbf{z}_{t}^{\mathrm{disc}}=\frac{1}{L-1}\cdot\mathrm{round}\left(\mathbf{z}_{t}^{\mathrm{cont}}\cdot(L-1)\right),$
@@
To generate symbolic ECG tokens, the quantized vector $\mathbf{z}_{t}^{\mathrm{disc}}$ is mapped to a unique token index using positional base conversion:
$\mathit{token\_id}_{t}=\sum_{i=1}^{D}v_{t,i}\cdot L^{i-1},$
where $v_{t,i}=\mathrm{round}\left(z_{t,i}^{\mathrm{disc}}\cdot(L-1)\right)\in\{0,1,\dots,L-1\}$. This yields $\mathit{token\_id}_{t}\in\{0,\dots,L^{D}-1\}$. These indices are used to look up embedding vectors from an ECG-specific vocabulary:
$\mathbf{Token}^{ECG}_{t}=\mathrm{ECG\_Vocabulary}[\mathit{token\_id}_{t}],$
where $\mathrm{ECG\_Vocabulary}\in\mathbb{R}^{|\mathcal{V}_{\mathrm{ECG}}|\times d_{LLM}}$ and $|\mathcal{V}_{\mathrm{ECG}}|=L^{D}$.
@@
$\hat{\mathbf{X}}=f_{\mathrm{dec}}(\mathbf{z}^{\prime}),\quad\mathcal{L}_{\mathrm{recon}}=\frac{1}{LT}\sum_{\ell=1}^{L}\sum_{t=1}^{T}\left(x_{\ell,t}-\hat{x}_{\ell,t}\right)^{2}.$
@@
The resulting ECG token sequence $\{\mathbf{Token}^{ECG}_{t}\}_{t=1}^{T^{\prime}}$ can be used as input to a language model. ECG tokens are initialized from the same distribution as the LLM’s text tokens. In the next stage, we leverage these ECG token sequences for autoregressive pretraining, using the same pretraining objective as for text tokens. Therefore, ECG token embeddings are drawn from the same embedding space as text tokens, the LLM can process ECG and text in a unified fashion.
@@
To integrate ECG signals into the language modeling framework, we pretrain the model on an autoregressive forecasting task over discretized ECG tokens. This allows the LLM to build temporal dependencies within ECG sequences and learn to embed symbolic ECG representations into its semantic space.
Given an input ECG signal $\mathbf{X}$, the tokenizer converts it into a discrete token sequence $\mathrm{Token}^{ECG}=\{\mathbf{t}_{1},\dots,\mathbf{t}_{T^{\prime}}\}\subseteq\mathrm{ECG\_Vocabulary}$, where each token is embedded in the same space as textual input. During pretraining, a random slice of this sequence is selected and split into a historical context and a prediction target (typically in a 9:1 ratio). The LLM is trained to autoregressively predict the next ECG tokens from the context:
$\mathcal{L}_{\mathrm{pretrain}}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{CE}\big{(}\mathrm{LM}(\mathrm{Token}^{ECG}_{\mathrm{input}}),\mathrm{Token}^{ECG}_{\mathrm{label}}\big{)},$
where $N$ is the batch size, and $\mathrm{LM}(\cdot)$ denotes the output logits of the LLM.
@@
In the final stage, we fine-tune the model on downstream clinical tasks, including ECG diagnostic report generation and open-ended question answering. This stage enables the pretrained model from last stage to align symbolic ECG representations with textual outputs in task-specific contexts. Following common instruction tuning protocols, we freeze the embedding table, output classification head, and entire LLM except for the linear projection layers (e.g., QKV and feedforward layers), which are adapted using LoRA.
Each training sample is structured as a text-style prompt that concatenates five components: (1) a brief dataset description (e.g., “The dataset contains electrocardiogram (ECG) time-series data.”), (2) a task-specific instruction (e.g., “Based on the given ECG signal embeddings, generate an ECG diagnostic report.”), (3) tabular features (including patient ID, age, sex, height, weight, recording date, nurse ID, device ID, and recording site), (4) an optional natural language question when the task is ECG-QA, and (5) ECG indicators (e.g., <|start_ecg|> and <|end_ecg|>) that mark the position for inserting ECG token sequences. The model is trained to autoregressively generate either a diagnostic report or an answer span, depending on the task. The supervision signal comes from expert-written reports or curated answers, enabling the model to learn both structured reasoning and domain-specific terminology. This unified format allows the model to flexibly handle both classification-type questions and free-form report generation, without needing task-specific heads or architectural changes.
@@
Ours TIMELLM TEST MEIT STMEM
W2V-
CMSC-
RLM
LLMTIME ChatTime
Fusion
Trans.
M3AE
MIMIC- IV-ECG
QA-Verify EM ACC 78.12 66.00 67.86 67.69 69.73 71.57 69.97 68.78 72.91
@@
MIMIC-IV-ECG
Report BLEU-1 44.99 14.49 28.40 27.20 24.89 35.89 28.92 26.29
@@
(a) PTB-XL Report
(b) PTB-XL QA-Verify
(c) PTB-XL QA-Choose
(d) PTB-XL QA-Query
Figure 3: Ablation study on the PTB-XL dataset. (a) Report generation evaluated by BLEU, METEOR, and ROUGE scores. (b–d) EM accuracy on QA-Verify, QA-Choose, and QA-Query tasks. We compare the full model (Ours) against three variants: (1) without discretization (w/o DISC), (2) without fine-tuning the LLM (w/o FT), and (3) without tabular features in the prompt (w/o TAB).
@@
We evaluate our method on two ECG-related language generation tasks: ECG-based question answering (ECG-QA) and diagnostic report generation (ECG-Report). The ECG-QA task follows the setup of (Tang et al. 2025), where 10,598 questions are curated based on six types of ECG attributes—extra systole, heart axis, signal noise, numerical indicators, SCP codes, and infarction stages—using recordings from PTB-XL (Wagner et al. 2020) and MIMIC-IV-ECG (Gow et al. 2023). Only single-ECG questions are considered. The ECG-Report task uses the same datasets with official diagnostic reports. Each dataset is split into training, validation, and test sets in a 7:1:2 ratio by patient ID to prevent data leakage. For MIMIC-IV-ECG, only ECGs without missing values are retained, and one ECG is randomly selected per patient, resulting in 78,358 ECGs. PTB-XL contains 21,797 ECGs in total.
@@
We compare our method with a range of TS-Text multimodal models. These include two TS-as-text baselines: ChatTime (Wang et al. 2025) and LLMTIME (Gruver et al. 2023), which directly feed raw numerical sequences into language models. We further consider five TS-embedding-based methods: TIME-LLM (Jin et al. 2024), TEST (Sun et al. 2024), MEIT (Wan et al. 2024), and LLaVA (Liu et al. 2023) with either ST-MEM (Na et al. 2024b) or W2V-SMSC-RLM (Oh et al. 2022) as the time-series encoder. In addition, we include two strong multimodal baselines specific to ECG-QA: Fusion Transformer (Oh et al. 2023) and M3AE (Chen et al. 2022b). Note that Fusion Transformer and M3AE do not support text generation and are thus excluded from ECG-Report evaluation. Moreover, M3AE requires pretraining with classification labels, which are not available in MIMIC-IV-ECG.

### v2

# HeartLLM: Discretized ECG Tokenization for LLM-Based Diagnostic Reasoning

@@ Abstract
Electrocardiography (ECG) plays a central role in cardiovascular diagnostics, yet existing automated approaches often struggle to generalize across clinical tasks and offer limited support for open-ended reasoning. We present HeartLLM, a novel framework that integrates time-series (TS) and language modeling by enabling large language models (LLMs) to process 12-lead ECG signals for clinical text generation tasks. Our approach discretizes continuous ECG embeddings into quantized codes using a lead-wise encoder and quantization module. These quantized codes are then mapped to an extended ECG vocabulary to form ECG tokens, enabling the model to process both ECG and natural language inputs within a unified framework. To bridge the modality gap, we pretrain the model on an autoregressive ECG token forecasting task, allowing the LLM to capture temporal dynamics through its inherent language modeling capability. Finally, we perform instruction tuning on both ECG question answering and diagnostic report generation. Without modifying the core model, HeartLLM achieves strong performance across tasks while maintaining generalization to out-of-distribution settings. Extensive experiments demonstrate the effectiveness of each component and highlight the potential of integrating discretized ECG tokens into LLMs for medical reasoning.
Codehttps://github.com/yangjinning/HeartLLM
@@ Introduction
Figure 1: UMAP visualizations of ECG and text embedding distributions across different models. (a-b) TEST and TIMELLM exhibit clear modality separation. (c) CLIP shows partially aligned clusters, but still fragmented modality boundaries. (d) HeartLLM aligns all ECG representations with text in a shared semantic space, indicating effective modality unification without explicit contrastive pairing.
@@
Cardiovascular diseases (CVDs) are the leading cause of mortality and morbidity worldwide (Di Cesare et al. 2023), and electrocardiography (ECG) remains a fundamental tool for early diagnosis and monitoring. With the increasing volume and complexity of ECG data in clinical settings, computer-aided interpretation has become an essential supplement to clinical expertise. Recent advances in deep learning have significantly improved performance in clinical decision support and ECG classification tasks (Ge et al. 2024; Xue et al. 2023a; Guhdar et al. 2025; Na et al. 2024b; Nejedly et al. 2021; Oh et al. 2022; Xue et al. 2023b), surpassing traditional signal-processing-based pipelines (Chen et al. 2022a). However, most existing models are tailored to closed-set classification and rely heavily on supervised training. When confronted with new diagnostic categories or unseen clinical tasks, they require extensive retraining with labeled data, limiting their scalability and adaptability. Moreover, these models are typically designed for narrow classification objectives and cannot support more flexible reasoning tasks such as open-ended question answering or narrative report generation.
@@
In parallel, large language models (LLMs) have demonstrated impressive generalization in vision-language domains (Alayrac et al. 2022; Huang et al. 2023), prompting efforts to extend such capabilities to physiological TS data like electrocardiograms (ECGs). However, this extension poses unique challenges: unlike symbolic and semantically structured text, ECG signals are continuous, noisy, and lack explicit semantic boundaries. These fundamental differences complicate the integration of ECG and text into a unified representation space. Recent studies have explored multimodal LLMs for clinical applications, including ECG-based question answering (Liu et al. 2024a; Zhao et al. 2024) and report generation (Wan et al. 2024). Most adopt a two-stage paradigm where an ECG encoder produces continuous representations that are subsequently consumed by a language model. However, this design suffers from two major limitations. First, the modality gap between continuous ECG embeddings and discrete textual semantics remains underexplored. Prior alignment strategies in time series–language modeling (Jin et al. 2024; Sun et al. 2024) or contrastive learning in vision–language settings (Radford et al. 2021) often fail to achieve consistent cross-modal alignment. These limitations, when directly applied to ECG, result in either overly clear or fragmented modality boundaries, as illustrated in Figure 1. Second, ECG signals exhibit high redundancy due to waveform repetition and stable low-variance segments. Directly feeding such continuous features into LLMs can lead to overfitting and reduced generalization to rare or subtle signal variations. These challenges highlight the need for a more structured and symbolic interface between ECG signals and language models.
@@
To address the challenges of bridging physiological signals and language models, we introduce HeartLLM, a unified framework that enables LLMs to reason over ECG signals in an open-ended, instruction-driven manner. Instead of relying on paired ECG–text supervision or complex cross-modal alignment losses, HeartLLM represents ECGs as discrete symbolic tokens, naturally compatible with the LLM’s vocabulary.
@@
Our method is built on three core contributions: First, we design a lead-wise encoder that processes each ECG lead independently using an identical encoder built from temporal inception enhancement blocks. This architecture captures waveform-specific temporal patterns while avoiding inter-lead interference, resulting in precise and interpretable representations. Second, we introduce a discretization-based ECG tokenizer that converts continuous encoder outputs into symbolic tokens via a fixed-scale quantizer (FSQ) (Mentzer et al. 2023). These tokens are optimized through autoregressive pretraining and embedded in a shared space with language, enabling seamless integration with LLMs without contrastive learning or paired ECG-text data. Third, we apply lightweight instruction tuning for ECG report generation and question answering. By updating only low-rank modules and prompting with structured clinical inputs, our method achieves strong zero-shot generalization across tasks and datasets with minimal supervision.
@@
HeartLLM is evaluated on two ECG understanding benchmarks: question answering and diagnostic report generation. It achieves state-of-the-art performance across multiple datasets and exhibits strong zero-shot generalization. Further analysis confirms the contribution of each design component, and visualizations illustrate interpretable attention over clinically meaningful waveform regions.
@@
Recent work has begun to explore the integration of LLMs into ECG interpretation tasks, particularly for open-ended question answering and diagnostic report generation. ECG-ReGen (Tang et al. 2025) adopts a retrieval-augmented framework that identifies similar historical cases to assist report generation in a zero-shot setting. PULSE (Liu et al. 2024a) constructs an instruction-tuning dataset based on ECG images, enabling the fine-tuning of image-text multimodal models. MEIT (Wan et al. 2024) introduces TS embeddings directly into the LLM’s attention layers to bridge modality differences, while (Tang et al. 2024) treats ECG features as soft prompts and applies meta-learning for rapid adaptation to new tasks. Despite these advances, most existing methods depend on paired ECG–text data to supervise modality alignment, which requires costly annotation and constrains applicability. In contrast, our approach aligns ECG and language representations through token-level pretraining, avoiding the need for paired supervision while maintaining flexibility across tasks.
@@
Large language models have been increasingly applied to TS tasks, typically via two main directions. One leverages LLMs to extract textual features to guide TS modeling (Wang et al. 2024a; Liu et al. 2025b, a). The other treats LLMs as core TS processors, mainly via two paradigms. The TS-as-text approach (Wang et al. 2025; Ansari et al. 2024; Wang et al. 2024b; Gruver et al. 2023) tokenizes numerical sequences into text for direct input into LLMs, but suffers from length and attention complexity limitations. The TS-embedding-based approach instead encodes signals into continuous embeddings and feeds them into frozen LLMs, often using prompt tuning (Zhou et al. 2023; Liu et al. 2024c). Some recent work aligns TS embeddings with language semantics, such as S2IP (Pan et al. 2024), TEST (Sun et al. 2024), and TIME-LLM (Jin et al. 2024), through contrastive or reprogramming strategies. Others like VITRO (Bellos et al. 2025) and UniTime (Liu et al. 2024b) propose task-specific vocabularies or reconstruction pretraining. However, most focus on forecasting or classification, with limited support for open-ended language tasks grounded in physiological signals. In contrast, our method discretizes ECG features into symbolic tokens, enabling autoregressive pretraining and unified ECG-text processing for clinical QA and report generation.
@@
Figure 2: Overview of the HeartLLM framework. The model consists of three stages: (1) ECG Tokenizer, where 12-lead ECG signals are encoded by 12 lead-wise encoders and discretized with FSQ into symbolic ECG tokens; (2) ECG Token Pretraining, where an LLM is autoregressively pretrained on ECG tokens to jointly optimize the extended ECG vocabulary and the model using teacher-forcing next-token prediction; and (3) QA/Report Tuning, where the pretrained model is adapted to downstream tasks using lightweight LoRA tuning. Prompts include structured tabular features and textual instructions to guide generation or question answering.
Given a collection of 12-lead ECG signals $X=\{x_{i}\}_{i=1}^{n}$, where each $x_{i}\in\mathbb{R}^{L\times T}$ consists of $L=12$ leads and $T=5000$ time steps , we consider two tasks: ECG question answering (ECG-QA) and diagnostic report generation (ECG-Report).
@@
In this task, the model generates a diagnostic report $r_{i}$ for each ECG signal $x_{i}$. The goal is to learn a mapping $f_{\mathrm{report}}:x_{i}\mapsto\hat{r}_{i}$ that produces clinically coherent and diagnostically relevant reports.
@@ Methods
We propose HeartLLM, a unified framework that enables large language models to perform open-ended clinical reasoning over ECG signals through symbolic representation and lightweight adaptation.
First, we design a lead-wise encoder that captures fine-grained temporal features independently for each ECG lead, enhancing representation fidelity without cross-lead interference. Second, we introduce a discretization mechanism that maps continuous features to symbolic tokens using an FSQ, forming an ECG-specific vocabulary compatible with LLMs. These tokens are pretrained via autoregressive prediction, naturally aligning time series and text without paired supervision. Finally, HeartLLM is instruction-tuned on clinical tasks such as ECG report generation and question answering, using only LoRA modules for efficient adaptation. An overview of the full framework is shown in Figure 2.
@@
Inception Block. The Inception Block applies multiple 1D convolutions with kernel sizes $\{1,3,5,7\}$ in parallel to extract multi-scale temporal features.
Temporal Inception Enhancement (TIE) Block. The outputs from the Inception Block are summed and passed through GroupNorm and a LeakyReLU activation, followed by a residual connection to preserve information flow:
@@
TIE blocks are applied sequentially to capture long-range signal dependencies, followed by convolutional downsampling layers that progressively compress the temporal resolution. This hierarchical structure allows the encoder to balance fine-grained waveform preservation with contextual abstraction. The encoded representations from all lead-wise encoders are concatenated along the lead dimension to form the final latent ECG representation $\mathbf{Z}\in\mathbb{R}^{T^{\prime}\times D^{\prime}}$.
@@
To discretize continuous ECG representations, each latent ECG representation at timestep $t$ $\mathbf{z}_{t}\in\mathbb{R}^{D^{\prime}}$ is projected into a lower-dimensional space $\mathbb{R}^{D}$ using a linear layer with parameters $\mathbf{W}\in\mathbb{R}^{D\times D^{\prime}}$ and $\mathbf{b}\in\mathbb{R}^{D}$:
@@
where $\sigma(\cdot)$ is the sigmoid function. The quantized code is obtained by uniformly quantizing each projected dimension into $K$ discrete levels:
$\mathbf{z}_{t}^{\mathrm{disc}}=\frac{1}{K-1}\cdot\mathrm{round}\left(\mathbf{z}_{t}^{\mathrm{cont}}\cdot(K-1)\right),$
@@
To generate symbolic ECG tokens, the quantized code $\mathbf{z}_{t}^{\mathrm{disc}}$ is mapped to a unique token index using positional base conversion:
$\mathit{token\_id}_{t}=\sum_{i=1}^{D}v_{t,i}\cdot K^{i-1},$
where $v_{t,i}=\mathrm{round}\left(z_{t,i}^{\mathrm{disc}}\cdot(K-1)\right)\in\{0,1,\dots,K-1\}$. This yields $\mathit{token\_id}_{t}\in\{0,\dots,K^{D}-1\}$. These indices are used to look up embedding vectors from an ECG-specific vocabulary:
$\mathrm{tok}^{ECG}_{t}=\mathrm{ECG\_Vocabulary}[\mathit{token\_id}_{t}],$
where $\mathrm{ECG\_Vocabulary}\in\mathbb{R}^{|\mathcal{V}_{\mathrm{ECG}}|\times d_{LLM}}$ and $|\mathcal{V}_{\mathrm{ECG}}|=K^{D}$.
@@
$\hat{X}=f_{\mathrm{dec}}(\mathbf{Z^{\prime}}),\quad\mathcal{L}_{\mathrm{recon}}=\frac{1}{LT}\sum_{\ell=1}^{L}\sum_{t=1}^{T}\left(x_{\ell,t}-\hat{x}_{\ell,t}\right)^{2}.$
@@
The resulting ECG token sequence $\{\mathrm{tok}^{ECG}_{t}\}_{t=1}^{T^{\prime}}$ can be used as input to a language model. ECG tokens are initialized from the same distribution as the LLM’s text tokens. In the next stage, we leverage these ECG token sequences for autoregressive pretraining, using the same pretraining objective as for text tokens. Therefore, ECG token embeddings are drawn from the same embedding space as text tokens, the LLM can process ECG and text in a unified fashion.
@@
To integrate ECG signals into the language modeling framework, we pretrain the model on an autoregressive forecasting task over discretized ECG tokens. This allows the LLM to build temporal dependencies within ECG sequences and learn to embed symbolic ECG tokens into its semantic space.
Given an input ECG signal $\mathbf{X}$, the tokenizer converts it into a discrete token sequence $\{\mathrm{tok}^{ECG}_{t}\}_{t=1}^{T^{\prime}}$, where each token is embedded in the same space as textual input. During pretraining, a random slice of this sequence is selected and split into a historical context and a prediction target (typically in a 9:1 ratio). The LLM is trained to autoregressively predict the next ECG tokens from the context:
$\mathcal{L}_{\mathrm{AR}}=\frac{1}{\sum_{i=1}^{N}|\mathcal{M}_{i}|}\sum_{i=1}^{N}\sum_{t\in\mathcal{M}_{i}}\mathrm{CE}\!\left(\mathrm{LM}_{\theta}(\mathrm{tok}_{i,\leq t}^{ECG}),\mathrm{tok}_{i,t+1}^{ECG}\right),$
where $N$ is the batch size, $\mathcal{M}_{i}$ is the set of prediction steps, $\mathrm{LM}_{\theta}(\cdot)$ is the LLM output logits over the ECG vocabulary, and $\mathrm{CE}(\cdot,\cdot)$ is the cross-entropy loss.
@@
In the final stage, we fine-tune the model on downstream clinical tasks, including ECG diagnostic report generation and open-ended question answering. This stage enables the pretrained model from the previous stage to align symbolic ECG representations with textual outputs in task-specific contexts. Following common instruction-tuning practice, the embedding table and output classification head are frozen, and only the linear projection layers of the LLM (e.g., QKV and feed-forward layers) are updated with LoRA.
Each training sample is structured as a prompt that concatenates five components: (1) a brief dataset description (e.g., “The dataset contains electrocardiogram (ECG) time-series data.”), (2) a task-specific instruction (e.g., “Based on the given ECG signal embeddings, generate an ECG diagnostic report.”), (3) tabular features (including patient ID, age, sex, height, weight, recording date, nurse ID, device ID, and recording site), (4) an optional natural language question when the task is ECG-QA, and (5) ECG indicators (e.g., <|start_ecg|> and <|end_ecg|>) that mark the position for inserting ECG token sequences. The model is trained to autoregressively generate either a diagnostic report or an answer span, depending on the task. The supervision signal comes from expert-written reports or curated answers, enabling the model to learn both structured reasoning and domain-specific terminology. This unified format allows the model to flexibly handle both classification-type questions and free-form report generation, without needing task-specific heads or architectural changes.
@@
Ours TIMELLM TEST MEIT STMEM W2V- CMSC- RLM LLMTIME ChatTime Fusion Trans. M3AE
MIMIC- IV-ECG QA-Verify EM ACC 78.12 66.00 67.86 67.69 69.73 71.57 69.97 68.78 72.91
@@
MIMIC-IV-ECG Report BLEU-1 44.99 14.49 28.40 27.20 24.89 35.89 28.92 26.29
@@
We evaluate our method on two ECG-related language generation tasks: ECG-based question answering (ECG-QA) and diagnostic report generation (ECG-Report). The ECG-QA task follows the setup of (Tang et al. 2025), where 10,598 questions are curated based on six types of ECG attributes—extra systole, heart axis, signal noise, numerical indicators, SCP codes, and infarction stages—using recordings from PTB-XL (Wagner et al. 2020) and MIMIC-IV-ECG (Gow et al. 2023). Only single-ECG questions are considered. The ECG-Report task uses the same datasets with official diagnostic reports. Each dataset is split into training, validation, and test sets in a 7:1:2 ratio by patient ID to prevent data leakage. For MIMIC-IV-ECG, only ECGs without missing values are retained, and one ECG is randomly selected per patient, resulting in 78,358 ECGs. PTB-XL contains 21,797 ECGs.
@@
We compare our method with a range of TS-Text multimodal models. These include two TS-as-text baselines: ChatTime (Wang et al. 2025) and LLMTIME (Gruver et al. 2023), which directly feed raw numerical sequences into language models. We further consider five TS-embedding-based methods: TIME-LLM (Jin et al. 2024), TEST (Sun et al. 2024), MEIT (Wan et al. 2024), and LLaVA (Liu et al. 2023) with either ST-MEM (Na et al. 2024a) or W2V-SMSC-RLM (Oh et al. 2022) as the ECG encoder. In addition, we include two strong multimodal baselines specific to ECG-QA: Fusion Transformer (Oh et al. 2023) and M3AE (Chen et al. 2022b). Note that Fusion Transformer and M3AE do not support text generation and are thus excluded from ECG-Report evaluation. Moreover, M3AE requires pretraining with classification labels, which are not available in MIMIC-IV-ECG.